supply are given by the equations:



Qd= 2000 – 400P + 0.5I +150Pb



Qs= -516 + 300P – 60W





where Qd and Qs is quantity of tourist e-book guides demanded and supplied, P is the price of a tourist e-book guide, I is household income, W is wage rate paid to tourist book guide laborers, and Pb is the price of a hardbound tourist book guide. Assume I is Php 2,300, W is Php10, and Pb is Php 21.40. Determine the equilibrium price and quantity of tourist e-book guides in this local market. Show your solution.

1
Expert's answer
2022-03-27T18:53:04-0400

"Let us take Q_{d}=Q_{s}"

"Q_{d}=2000-400p+0.5\times I+150\times pb"

"Q_{d}= 2000-400p + 0.5\times2300 + 150\times21.40"

"Q_{d}= 6360-400p"


"Q_{s} =-516+300p -60\times W"

"Q_{s} = -516+300p -60\times10"

"Q_{s} = - 516+300p - 600"

"Q_{s} = -1116+300p"


"isolate the variable P"

"Since, Q_{d}=Q_{s};"

"6360-400p= -1116+300p"

"6360+1116=400p+300p"

"7476=700p"

"\frac{7476}{700p}=10.68"

"p=10.68"

"equilibrium price =10.68"


quantity of tourist e-book guides in this local market

"6360-400\times p"

"6360-400\times10.68"

"6360-4272"

"Quantity= 2088"